OpenStudy (anonymous):

How is solving a literal equation different from solving a one- or two-step equation? plz

OpenStudy (anonymous):

plz help me this is my last question @ikram002p

OpenStudy (anonymous):

@satellite73 plz help

OpenStudy (anonymous):

@kropot72 plz help me

OpenStudy (ikram002p):

more step and justification each step in literal equation

OpenStudy (ikram002p):

one or two i think they are sometimes linear but i cant see clear difference for what you want

OpenStudy (ikram002p):

let me think mmm

OpenStudy (anonymous):

huh? i dont really know i just copied and pasted the question

OpenStudy (ikram002p):

mmm @ganeshie8 do you have an idea ?

OpenStudy (anonymous):

plz its 1am and i need to finish i will take any of your advice

OpenStudy (ikram002p):

ok if we have this equation 2x+3=1 and wanna solve for x literal equation first we add -3 for both sides 2x+3-3=1-3 2x+0=-2 2x=-2 second step multiply with 1/2 both sides 2x * 1/2 = -2 * 1/2 1 x = -1 x=1 the same thing with a one- or two-step equation? 2x+3=1 2x=1-3=-2 then devide by 2 x=-1 mmm thats it :-\ to me both are the same

OpenStudy (ikram002p):

will this is what i see in literal equation there is more steps and justification for each step